Find the zeros of x^3 - x^2 + 4x - 4
Thank you!

Answer by jim_thompson5910(35256)

x^3 - x^2 + 4x - 4 = 0

(x^3 - x^2) + (4x - 4) = 0

x^2(x - 1) + (4x - 4) = 0

x^2(x - 1) + 4(x - 1) = 0

(x^2 + 4)(x - 1) = 0

x^2 + 4 = 0 or x - 1 = 0

x^2 = -4 or x = 1

x = sqrt(-4), x = -sqrt(-4) or x = 1

x = 2i, x = -2i or x = 1

where i = sqrt(-1)